# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v2

@@ Abstract
Aerial remote sensing enables efficient large-area surveying, but accurate direct object-level measurement remains difficult in complex natural scenes. Recent advancements in 3D vision, particularly learned radiance-field representations such as NeRF and 3D Gaussian Splatting, have begun to raise the ceiling on reconstruction fidelity and densifiable geometry from posed imagery. Nevertheless, direct aerial measurement of important natural attributes such as tree diameter at breast height (DBH) remains challenging. Trunks in aerial forest scans are distant and sparsely observed in image views: at typical operating altitudes, stems may span only a few pixels. With these constraints, conventional reconstruction methods leave breast-height trunk geometry weakly constrained. We present TreeDGS, an aerial image reconstruction method that leverages 3D Gaussian Splatting as a continuous, densifiable scene representation for trunk measurement. After SfM–MVS initialization and Gaussian optimization, we extract a dense point set from the Gaussian field using RaDe-GS’s depth-aware cumulative-opacity integration and associate each sample with a multi-view opacity reliability score. Then, we estimate DBH from trunk-isolated points using opacity-weighted solid-circle fitting. Evaluated on 10 plots with field-measured DBH, TreeDGS reaches 4.79,cm RMSE (about 2.6 pixels at this GSD) and outperforms a state-of-the-art LiDAR baseline (7.91,cm RMSE). This shows that TreeDGS can enable accurate, low-cost aerial DBH measurement (Fig. LABEL:fig1).
@@ 1 Introduction
Aerial remote sensing has quickly become a cornerstone of modern environmental monitoring for its ability to survey large areas rapidly and cost efficiently [1, 7, 36]. Yet, despite major gains in sensor resolution, platform stability, and reconstruction pipelines, extracting direct, object-level measurements from aerial imagery remains difficult in structurally complex natural scenes [8, 15]. Forested environments are challenging: heterogeneous geometry, self-similar textures, frequent occlusion, and strong appearance variability across lighting conditions all weaken the visual cues that aerial reconstruction methods rely upon [15].
@@
A prominent example is tree diameter at breast height (DBH)—a foundational forestry variable that aerial remote sensing struggles to recover reliably. Its importance for forestry cannot be understated: DBH measures an individual tree’s size and growth stage, and is a primary input to standard allometric models and reporting workflows that estimate wood volume, biomass, and carbon [2, 17, 5, 18]. In turn, these allometric models underpin decision making used across forestry operations and policy, ranging from timber inventory and silvicultural planning to carbon offset quantification, wildfire risk mitigation, and long-term ecosystem management [41, 3].
@@
DBH estimation from above-canopy aerial sensing is fundamentally constrained by observability. In typical UAV image surveys flown at operational altitudes, stems are not “large objects.” In our data, captured at $\sim$70 m above ground ($\mathrm{GSD}\approx 1.84$ cm/px), stems in the evaluation plots span only $\sim$15 pixels across per view on average (Tab. 1), and can be just a few pixels wide for smaller trees (Fig. 2). This pixel scarcity is compounded by view scarcity: the breast-height band ($h_{\mathrm{BH}}\approx 1.37$ m in our field protocol) is frequently occluded by crowns, branches, and understory vegetation, so each stem may be seen only in a handful of oblique glimpses through noisy canopy gaps. Therefore, the core geometric challenge of measurement is recovering enough faithful breast-height trunk surface to support stable circle or cylinder fitting under a regime of few pixels, few views, and partial visibility.
Active sensing can mitigate some of these constraints by providing explicit range measurements rather than relying on image texture. Terrestrial laser scanning (TLS), for example, is well known to support robust DBH measurements because it can produce dense, well-populated stem cross-sections [30, 38, 20]. However, the mostly downward-looking trajectories of airborne or UAV laser scanning do not typically yield equally dense returns on the trunk at breast height: even with high point density, trunk hits near $h_{\mathrm{BH}}$ can be sparse, uneven, and contaminated by surrounding vegetation, complicating per-tree stem isolation and reliable fitting in cluttered stands [25, 35, 24]. These limitations help explain why aerial LiDAR has seen broader operational adoption for canopy structure and terrain products rather than for direct DBH measurement [13, 31, 39, 47, 6].
RGB-only UAV photogrammetry has nevertheless made substantial progress for other forest attributes, including canopy height models and structural mapping [8, 15, 45, 44]. But DBH remains difficult from aerial imagery alone, even under careful survey design with aggressive overlap and strong oblique views [34]. In the pixel- and view-limited regime described above, conventional SfM–MVS pipelines often fail to recover a coherent breast-height surface. Sparse SfM relies on repeatable keypoints and long feature tracks [40], but distant trunks provide weak texture and are frequently interrupted by foliage occlusions, yielding unstable or missing correspondences. View-dependent MVS densification [9] then tends to drop out on trunks when visibility is intermittent and texture cues are weak.
@@
Recent radiance-field representations, such as Neural Radiance Fields (NeRF) [33] and 3D Gaussian Splatting (3DGS) [22], can improve reconstruction fidelity from posed imagery and offer a promising path forward for thin structures that are only weakly observed in individual frames. Early forestry applications have begun exploring radiance fields for forest monitoring and tree-scale reconstruction [14, 23, 42, 27], but the problem of direct DBH measurement from stand-off, above-canopy UAV surveys remains largely unaddressed. Moreover, translating a “fluffy” radiance-field reconstruction into a reliable diameter requires a measurement-oriented extraction step: a sufficiently dense sampling at breast height and a way to prioritize samples that are consistently supported across views. Most forestry applications of radiance fields to date have focused on close-range terrestrial [27] imagery which have a different set of constraints, leaving aerial DBH measurement largely unaddressed by radiance field methods.
@@
We introduce TreeDGS: an aerial reconstruction pipeline that uses 3D Gaussian Splatting as a measurement-oriented scene representation for estimating DBH from commodity UAV RGB imagery under pixel- and view-limited trunk observations. Starting from standard SfM camera poses and an MVS initialization, we optimize a RaDe-GS model [50] to obtain a continuous Gaussian field whose covariances and opacities are refined by multi-view photometric consistency. Then, we extract a dense point set using depth-aware cumulative-opacity integration, retaining samples whose accumulated opacity indicates a reliably occupied surface under a fixed threshold. To mitigate spurious samples from foliage and partial occlusions, we compute a per-point multi-view opacity support score as a reliability signal and retain trunk-isolated points via 3D semantic segmentation. Finally, we estimate DBH with opacity-weighted, slice-wise circle fitting, yielding diameter estimates even when the trunk is only weakly observed in individual images.
Our contributions are summarized as follows:
  • We present TreeDGS, a method that directly reconstructs and measures DBH from above-canopy UAV RGB imagery by using a 3D Gaussian Splatting scene representation in a pixel-limited regime.
  • We introduce a depth- and opacity-aware point extraction procedure based on RaDe-GS that densifies the Gaussian field into a surface-consistent point set and attaches a multi-view opacity support score, enabling reliability-weighted geometric fitting.
  • We demonstrate real-world, field-validated performance in a managed loblolly pine forest with dense understory and varied growth patterns (10 plots), achieving 4.79 cm RMSE ($\sim$2.6 pixels at our GSD) against ground-truth tape DBH and outperforming an ultra-high-resolution UAV LiDAR baseline by 39.4% in RMSE.
@@ 2 Materials
Data were collected in a managed Pinus taeda (loblolly pine) stand in southeastern Arkansas, USA. We established ten 0.2-acre circular plots (plots 1–10; radius 16.05 m; Fig. 3). At each plot center, we marked a circular plot boundary with a 16.05 m radius and measured all stems within the boundary. Each included tree was assigned a unique ID. For each tree, we recorded: (i) bearing and distance from the plot center; (ii) DBH measured with a diameter tape at 1.37 m above ground level; and (iii) height measured using a laser rangefinder.
We surveyed plot centers using ORS/NTRIP network corrections. For each plot, a survey-grade GNSS rover mounted on a survey pole and stabilized with a bipod was centered over the plot and occupied for 30 minutes in static mode while receiving RTCM corrections via an NTRIP caster. Plot-center coordinates were derived using a differential GNSS workflow, following forest-canopy positioning best-practice recommendations described by Strunk et al. [43]. Corrected plot-center coordinates were then exported for co-registration with the aerial products.
Figure 4: RGB and LiDAR acquisition patterns. (a) RGB imagery was collected with high overlap and mixed viewing angles to support SfM/MVS and improve trunk visibility. (b) LiDAR was flown with dense grid (lawnmower) flight lines to obtain uniform coverage across the plot network.
Table 1: Per-plot summary for the manually paired subset used in per-tree evaluation. #Trees denotes the number of stems in each plot with a verified field-to-reconstruction association; the All row reports pooled statistics across all plots.
[…]
@@
UAV RGB imagery.
For the full set of plots, we collected overlapping RGB imagery using a DJI Matrice 4E in two campaigns. Imagery was captured with high overlap and mixed viewing geometry (including oblique views at 70 and 90 m altitude) to support SfM/MVS reconstruction and increase the chance of observing trunk surfaces through gaps in the canopy and understory. To improve absolute positioning consistency across flights, we corrected the onboard UAV GNSS trajectory via post-processed kinematic (PPK) processing using reference data from a nearby CORS station (ARMO), and used the corrected solution to update image geotags before reconstruction. An example UAV RGB flight trajectory is shown in Figure 4(a).
Figure 5: TreeDGS pipeline. RGB images are reconstructed with SfM/MVS and represented as 3D Gaussian splats $G$. We surface-sample dense points using opacity $\alpha$ and weights $w$, segment trunks $T$, and estimate DBH by opacity-weighted circle fitting on trunk slices.
UAV LiDAR
@@
Manual field-to-reconstruction association.
Across all plots, we measured 458 stems in the field. We manually paired 210 pine stems (the target species in this study) with reconstructed trunk instances; Table 1 reports per-plot pine stem counts and DBH distribution statistics.
@@ 3.1 Problem Statement
Our approach is geometry-driven: we reconstruct camera poses and an initial 3D structure with SfM, densify it with OpenMVS [4], optimize a Gaussian Splatting [22, 50] scene representation, and then measure DBH from trunk-only geometry using opacity-aware sampling and robust fitting. Fig. 5 overviews the workflow; we summarize it as:
@@
where $\mathbf{P}_{i}$ are the calibrated camera matrices, $\mathcal{X}_{\text{dense}}$ is a densified point set from OpenMVS, $\mathcal{G}$ is the optimized Gaussian field, and $\mathcal{T}_{t}$ denotes trunk points for tree instance $t$.
@@ 3.2 Structure-from-Motion and Multi-View Stereo
Coarse-to-fine matching.
Because our UAV imagery is high resolution, we perform matching in a coarse-to-fine manner [26]: we first obtain coarse correspondences on downsampled images to establish robust global alignment, then refine matches locally at higher resolution. This reduces compute while maintaining accurate pixel localization needed for stable pose estimation.
SfM reconstruction.
SfM takes the matched correspondences across overlapping image pairs and estimates: (i) camera poses $\{\mathbf{P}_{i}\}$ and (ii) a sparse 3D point set. We use standard robust estimation and bundle adjustment from GLOMAP [37] to refine the reconstruction. We denote the resulting calibrated cameras as:
@@
These cameras are used both for Gaussian training (Sec. 3.3) and for the multi-view opacity tests during surface sampling (Sec. 3.4).
Coarse densification with OpenMVS [4].
After SfM, we densify geometry using OpenMVS. OpenMVS produces depth information and a denser point set $\mathcal{X}_{\text{dense}}$, which provides stronger geometric support than the sparse SfM points in trunk regions. In our pipeline, $\mathcal{X}_{\text{dense}}$ serves as an initialization / prior for Gaussian optimization in Section 3.3.
Figure 6: Comparison of back-projection points with multi-view depth consistency and the proposed sample points. (a) Depth-fusion baseline: rendered depth points are back-projected into 3D and kept only if they pass multi-view depth-consistency checks. This produces large sheet-like artifacts near the ground/understory and incomplete stem surfaces (red dashed boxes), despite appearing consistent across views. (b) Our opacity-guided surface sampling with depth-aware point-wise compositing gives a stem geometry suitable for DBH estimation.

### v3

@@ Abstract
Aerial remote sensing efficiently surveys large areas, but accurate direct object-level measurement remains difficult in complex natural scenes. Advancements in 3D computer vision, particularly radiance field representations such as NeRF and 3D Gaussian splatting, can improve reconstruction fidelity from posed imagery. Nevertheless, direct aerial measurement of important attributes like tree diameter at breast height (DBH) remains challenging. Trunks in aerial forest scans are distant and sparsely observed in image views; at typical operating altitudes, stems may span only a few pixels. With these constraints, conventional reconstruction methods have inaccurate breast-height trunk geometry. TreeDGS is an aerial image reconstruction method that uses 3D Gaussian splatting as a continuous scene representation for trunk measurement. After SfM–MVS initialization and Gaussian optimization, we extract a dense point set from the Gaussian field using RaDe-GS’s depth-aware cumulative-opacity integration and associate each sample with a multi-view opacity reliability score. Then, we isolate trunk points and estimate DBH using opacity-weighted solid-circle fitting. Evaluated on 10 plots with field-measured DBH, TreeDGS reaches 4.79 cm RMSE (about 2.6 pixels at this GSD) and outperforms a LiDAR baseline (7.66 cm RMSE). This shows that TreeDGS can enable accurate, low-cost aerial DBH measurement (Figure LABEL:fig1).
@@ 1 Introduction
Aerial remote sensing has quickly become a cornerstone of modern environmental monitoring for its ability to survey large areas rapidly and cost efficiently  [1, 7, 37]. Yet, despite major gains in sensor resolution, platform stability, and reconstruction pipelines, extracting direct, object-level measurements from aerial imagery remains difficult in structurally complex natural scenes [8, 15]. Forested environments are particularly challenging as heterogeneous geometry, self-similar textures, frequent occlusion, and strong appearance variability across lighting conditions all weaken the visual cues that aerial reconstruction methods rely on [15].
@@
A prominent example is tree diameter at breast height (DBH). This describes an individual tree’s size and growth stage. DBH is an important forestry variable that is a primary input to standard allometric models and reporting workflows to estimate wood volume, biomass, and carbon [2, 17, 5, 18]. In turn, these allometric models underpin decision making used across forestry operations and policy, ranging from timber inventory and silvicultural planning to carbon offset quantification, wildfire risk mitigation, and long-term ecosystem management [42, 3].
@@
DBH estimation from above-canopy aerial sensing is fundamentally limited by observability. In typical UAV image surveys flown at operational altitudes, stems are not large objects. In our data, captured at $\sim$70 m above ground ($\mathrm{GSD}\approx 1.84$ cm/px), stems in the evaluation plots span only $\sim$15 pixels across per view on average (Table 1), and they can be just a few pixels wide for smaller trees (Figure 2). This pixel scarcity is compounded by view scarcity: the breast-height band ($h_{\mathrm{BH}}\approx 1.37$ m in our field protocol) is frequently occluded by crowns, branches, and understory vegetation, so each stem may be seen only in a handful of oblique views through noisy canopy gaps. Thus, the core geometric challenge of measurement is to recover enough faithful breast-height trunk surface samples to support stable circle or cylinder fitting.
@@
Active sensing can mitigate some of these limitations by providing explicit range measurements rather than relying on image texture. Terrestrial laser scanning (TLS), for example, is well known to support robust DBH measurements because it can produce dense, well-populated stem cross-sections [31, 39, 21]. However, the mostly downward-looking trajectories of airborne or UAV laser scanning (ULS) do not typically yield equally dense returns on the trunk at breast height: even with high point density, trunk hits near $h_{\mathrm{BH}}$ can be sparse, uneven, and contaminated by surrounding vegetation, complicating per-tree stem isolation and reliable fitting in cluttered stands [26, 36, 25]. These limitations help explain why UAV LiDAR has seen broader operational adoption for canopy structure and terrain products than for direct DBH measurement [13, 32, 40, 49, 6].
Nevertheless, RGB-only UAV photogrammetry has made substantial progress for other forest attributes, including canopy height models and structural mapping [8, 15, 46, 45]. In contrast, DBH remains difficult from aerial imagery alone, even under careful survey design with aggressive overlap and strong oblique views [35]. In the pixel- and view-limited regime described above, conventional SfM–MVS pipelines often yield incomplete and fragmented breast-height trunk surfaces that are insufficient for stable diameter fitting. Sparse SfM relies on repeatable keypoints and long feature tracks [41], but distant trunks provide weak texture and are frequently interrupted by foliage occlusions, yielding unstable or missing correspondences. View-dependent MVS densification [9] then tends to drop out on trunks when visibility is intermittent and texture cues are weak. Still, the partial geometry recovered by MVS can serve as a useful coarse prior for later optimization.
@@
Recent radiance field representations, such as Neural Radiance Fields (NeRF) [34] and 3D Gaussian Splatting (3DGS) [23], can improve reconstruction fidelity from posed imagery and offer a promising path forward for thin structures that are only weakly observed in individual frames. Early forestry applications have begun exploring radiance fields for close-range terrestrial forest monitoring and tree-scale reconstruction [14, 24, 43, 28], but the problem of direct DBH measurement from stand-off, above-canopy UAV surveys remains unsolved. Moreover, translating a “fluffy” radiance field reconstruction into a reliable diameter measurement requires an extraction step, such that we can densely sample trees at breast height and prioritize samples that are consistently supported across views.
@@
We introduce a method—TreeDGS—that directly measures DBH from above-canopy UAV RGB imagery from a 3D Gaussian splatting scene reconstruction under pixel- and view-limited trunk observations. Starting from standard SfM camera poses and an MVS initialization, we optimize a RaDe-GS model [52] to obtain a continuous Gaussian field whose covariances and opacities are refined by multi-view photometric consistency. Then, we extract a dense point set using depth-aware cumulative-opacity integration, retaining samples whose accumulated opacity indicates a reliably occupied surface under a fixed threshold. To mitigate spurious samples from foliage and partial occlusions, we compute a per-point multi-view opacity support score as a reliability signal and retain trunk-isolated points via 3D semantic segmentation. Finally, we estimate DBH with opacity-weighted, slice-wise circle fitting, yielding stable diameter estimates even when the trunk is only weakly observed in individual images. We demonstrate real-world, field-validated performance in a managed loblolly pine forest with dense understory and varied growth patterns (10 plots), achieving 4.79 cm RMSE ($\sim$2.6 pixels at our GSD) against ground-truth tape DBH and outperforming an ultra-high-resolution UAV LiDAR baseline by 37.5% in RMSE.
@@ 2 Materials
Table 1: Per-plot summary for the manually paired subset used in per-tree evaluation. #Trees denotes the number of stems in each plot with a verified field-to-reconstruction association; the “All” row reports pooled statistics across all plots.
Plot #Trees DBH Mean $\pm$ SD (cm) DBH Range (cm)
@@
We collected data in a managed Pinus taeda (loblolly pine) stand in southeastern Arkansas, USA. We established ten 0.2-acre circular plots (plots 1–10; radius 16.05 m; Figure 3). At each plot center, we marked a circular plot boundary with a 16.05 m radius and measured all stems within the boundary. Then, we assigned a unique ID to each included tree. For each tree, we recorded the following: (i) Bearing and distance from the plot center; (ii) DBH measured with a diameter tape at 1.37 m above ground level; and (iii) height measured using a laser rangefinder.
We surveyed plot centers using ORS/NTRIP network corrections. For each plot, we centered a survey-grade GNSS rover mounted on a survey pole and stabilized with a bipod over the plot and occupied for 30 min in static mode while receiving RTCM corrections via an NTRIP caster. Then, we derived plot-center coordinates using a differential GNSS workflow, following forest-canopy positioning best-practice recommendations described by Strunk et al. [44]. Finally, we exported corrected plot-center coordinates for co-registration with the aerial products.
UAV RGB imagery. For the full set of plots, we collected overlapping RGB imagery using a DJI Matrice 4E (SZ DJI Technology Co., Ltd., Shenzhen, China) in two campaigns. We captured imagery with high overlap and mixed viewing geometry (including oblique views at 70 and 90 m altitude) to support SfM/MVS reconstruction and increase the chance of observing trunk surfaces through gaps in the canopy and understory. To improve absolute positioning consistency across flights, we corrected the onboard UAV GNSS trajectory via post-processed kinematic (PPK) processing using reference data from a nearby CORS station (ARMO), and we used the corrected solution to update image geotags before reconstruction. An example UAV RGB flight trajectory is shown in Figure 4a.
Figure 4: RGB and LiDAR acquisition patterns. (a) We collected RGB imagery with high overlap and mixed viewing angles to support SfM/MVS and improve trunk visibility. (b) We captured LiDAR with dense grid (lawnmower) flight lines to obtain uniform coverage across the plot network. Red polylines indicate UAV flight trajectories; in (a) points are colored by RGB, while in (b) LiDAR points are colored by height above ground (blue low to yellow high).
@@
The resulting georeferenced point clouds serve as a baseline geometry source for trunk isolation and DBH fitting.
Pairing workflow and quality control. Across all plots, we measured 458 stems in the field. To create the per-tree evaluation set, we paired each field-inventoried pine stem to a reconstructed trunk instance using a standardized, spatially driven workflow. For each plot, we converted the field-measured (bearing and distance) from the surveyed plot center into georeferenced $(x,y)$ coordinates. For the reconstructed data (TreeDGS and LiDAR), we represented each trunk instance by the centroid of its segmented trunk points near breast height. First, we generated candidate associations using nearest-neighbor matching within the plot boundary, and then we manually verified each pair in a plot-level visualization that overlays the field points, reconstructed instances, and high-resolution aerial context (orthomosaic/point cloud; Figure 5). Two authors performed pairing; uncertain or ambiguous cases (e.g., clustered stems, partially reconstructed trunks, or missing instances) were discussed and, if unresolved, excluded from the paired set. This resulted in 210 pair stems. During fieldwork, we also recorded a plot walkthrough video, which served as an additional reference to confirm tree neighborhoods when needed. Table 1 reports the per-plot pine stem counts and DBH distribution statistics.
Figure 5: Example of field-to-reconstruction stem matching. Plot-level overlay illustrating the manual pairing between field-inventoried pine stems and reconstructed trunk instances (after removing small non-pine trees). Lines connect matched pairs. The mean planimetric offset is 0.776 m for field-to-TreeDGS matches (210 pairs across 10 plots) and 0.270 m for LiDAR-to-TreeDGS matches. These offsets reflect combined field distance/bearing uncertainty and georegistration error, and they are used only as a sanity check after visual verification.
@@ 3.1 Problem Statement
Our approach is geometry-driven: we reconstruct camera poses and an initial 3D structure with SfM, densify it with OpenMVS [4], optimize a Gaussian splatting [23, 52] scene representation, and then measure DBH from trunk-only geometry using opacity-aware sampling and robust fitting (Figure 6):
@@
where $\mathbf{P}_{i}$ are the calibrated camera matrices, $\mathcal{X}_{\text{dense}}$ is a densified point set from OpenMVS, $\mathcal{G}$ is the optimized Gaussian field, and $\mathcal{T}_{t}$ denotes trunk points for tree instance $t$. Sections 3.23.6 detail the SfM/OpenMVS reconstruction, Gaussian optimization, trunk extraction through 3D segmentation [50] with opacity cues, and the final DBH fitting procedure.
Figure 6: TreeDGS pipeline. RGB images are reconstructed with SfM/MVS and represented as 3D Gaussian splats $G$. We surface-sample dense points using opacity $\alpha$ and weights $w$, segment trunks $T$, and estimate DBH by opacity-weighted circle fitting on multiple trunk slices. Inset colors: $\alpha$ and $w$ are visualized with a low-to-high colormap (cool colors = low values; warm colors = high values); in $T$, trunk points are red and non-trunk points are green/yellow.
@@ 3.2 Structure-from-Motion and Multi-View Stereo
Coarse-to-fine matching. Because our UAV imagery is high resolution, we perform matching in a coarse-to-fine manner [27]: we first obtain coarse correspondences on downsampled images to establish robust global alignment, and we then refine matches locally at higher resolution. This reduces computation while maintaining accurate pixel localization needed for stable pose estimation.
SfM reconstruction. SfM takes the matched correspondences across overlapping image pairs and estimates the following: (i) camera poses $\{\mathbf{P}_{i}\}$ and (ii) a sparse 3D point set. We use standard robust estimation and bundle adjustment from GLOMAP [38] to refine the reconstruction. We keep notation minimal and denote the resulting calibrated cameras as:
@@
These cameras are used both for Gaussian training (Section 3.3) and for the multi-view opacity tests during surface sampling (Section 3.4).
Coarse densification with OpenMVS [4]. After SfM, we densify geometry using OpenMVS. OpenMVS produces depth information and a denser point set $\mathcal{X}_{\text{dense}}$, which provides stronger geometric support than sparse SfM points in many well-observed regions (e.g., ground, canopy, and partially visible trunk segments). We note that $\mathcal{X}_{\text{dense}}$ is not assumed to be complete or coherent in the breast-height band. In our setting, MVS trunk coverage can be sparse and fragmented due to occlusion and weak texture. Instead, we use $\mathcal{X}_{\text{dense}}$ only as a coarse prior to initialize the Gaussian centers for faster and more stable optimization, following common 3DGS practice [23, 19, 47].